# From XAI to Stories: A Factorial Study of LLM-Generated Explanation Quality

### v1

@@ Abstract
Fabian Lukassen1, Jan Herrmann2, Christoph Weisser1, Benjamin Saefken1,3, Thomas Kneib1 1University of Göttingen  2BASF SE  3TU Clausthal fabian.lukassen@stud.uni-goettingen.de, jan.herrmann@basf.com, christoph.weisser@oxon.org benjamin.saefken@tu-clausthal.de, tkneib@uni-goettingen.de
@@ 1 Introduction
The deployment of ML models in critical domains such as healthcare, finance, and energy management has intensified the need for model interpretability (guidotti2018survey). XAI methods like SHAP (lundberg2017unified) and LIME (ribeiro2016why) have emerged as essential tools, providing quantitative feature attributions that reveal which inputs drive model predictions. However, these technical outputs – typically numerical importance scores or attribution values – remain inaccessible to non-expert users who lack the statistical background to interpret them (zytek2024llms).
@@
Recent work has demonstrated that LLMs can transform XAI output into NLEs that non-experts can better understand (Figure 1), with user studies validating this approach (zytek2024explingo; cedro2024graphxain).
@@
However, previous work on transforming XAI outputs into NLEs does not systematically study the effects of ML model choice, XAI method, LLM, and prompting strategy on NLE quality. We address this gap with a factorial study grounded in Design of Experiments (DOE) principles (collins2014factorial), focusing on time-series forecasting. A domain where temporal dependencies and seasonal patterns make XAI outputs especially challenging to interpret. The factorial combination of these four factors (see Figure 2 for their levels) yields 660 NLEs, which we evaluate using G-Eval (liu2023geval), an LLM-as-judge framework with dual judges and four evaluation criteria. Our main contributions are:
@@ 2.1 Explainable AI Methods
Among post-hoc explanation methods, SHAP and LIME have emerged as the two most widely adopted techniques for interpreting black-box models (garreau2020explaining).
SHAP (lundberg2017unified) provides theoretically grounded attributions based on Shapley values from cooperative game theory. For a prediction $f(x)$ with feature set $\mathcal{F}$, SHAP assigns each feature $i$ an importance value $\phi_{i}$ computed as a weighted average of marginal contributions across all possible feature coalitions $S$:
@@
LIME (ribeiro2016why) takes a different approach, fitting interpretable local surrogate models. Given instance $x$ and black-box model $f$, LIME learns a locally faithful linear model $g$ by solving:
@@
Both methods have been widely adopted (guidotti2018survey), applied in complex machine learning pipelines (semnani2024amachine; semnani2025from), and extended to specialized variants including TreeSHAP (lundberg2020local). However, despite their mathematical rigor, XAI outputs require statistical expertise to interpret, particularly for time-series data where temporal dependencies add complexity (theissler2022explainable). zytek2024llms identify this as a fundamental barrier, noting that “feature importance scores are only useful if users can understand what they mean.”
In contrast, classical statistical models like SARIMAX offer inherent interpretability through coefficients and confidence intervals (box2015time).
@@ 2.2 Natural Language Explanations from XAI
The transformation of XAI outputs to natural language has progressed from template-based systems to LLM-based approaches. martens2024xaistories introduced XAIstories, generating narratives from SHAP values, with user studies showing participants found them convincing and useful for communicating with laypeople. zytek2024explingo developed Explingo, also using SHAP, combining XAI-to-narrative transformation with an LLM-based grader. cedro2024graphxain extended this to graph neural networks using GNNExplainer. For time-series specifically, shajalal2024forecastexplainer used DeepLIFT for LSTM-based energy forecasting, while aksu2024xforecast proposed XForecast with simulatability metrics, finding that an LLM’s numerical reasoning ability matters more than model size for accurate explanations. Domain applications include chemistry (wellawatte2023extracting).
@@ 2.3 LLM Prompting Strategies
The effectiveness of LLMs depends heavily on prompting. Foundational techniques include zero-shot and few-shot in-context learning (brown2020language), while more advanced strategies target reasoning and reliability. Chain-of-thought (CoT) prompting elicits step-by-step reasoning (wei2022chain), which self-consistency extends by sampling multiple reasoning paths and selecting the most common answer (wang2023selfconsistency). Other approaches include reflexion, where models iteratively critique and refine their outputs (shinn2023reflexion); meta-prompting, which uses structured templates for multi-step reasoning (suzgun2024metaprompting); and role-based prompting, which assigns personas to guide generation (kong2024betterzeroshotreasoning).
@@ 2.4 LLM-as-Judge Evaluation
Traditional metrics (BLEU, ROUGE) correlate poorly with human judgment for open-ended text generation (liu2023geval). The LLM-as-judge paradigm has emerged as a scalable alternative (zheng2023judging), with three main evaluation paradigms (gu2024llmasjudge): pointwise (scoring individual outputs independently), pairwise (comparing two outputs to select the better one), and listwise (ranking multiple outputs simultaneously).
liu2023geval introduced G-Eval, a pointwise framework combining chain-of-thought prompting with probability-weighted scoring. G-Eval first generates evaluation steps from a criterion (e.g., “coherence”), then uses these steps to assess outputs via a form-filling paradigm. To obtain fine-grained scores, G-Eval computes:
@@
However, LLM judges exhibit systematic biases (gu2024llmasjudge): positional bias (preference for first or last positions in ranking tasks), self-preference bias (favoring outputs from the same LLM family), and verbosity bias (favoring longer responses). Mitigation strategies include randomizing presentation order, using judges from different model families than generators, and supplementing absolute scores with ranking-based evaluation.
@@ 3.2 Time-Series Forecasting Task
We use the UCI Individual Household Electric Power Consumption dataset (hebrail2012individual), containing 2,075,259 minute-level measurements of a French household’s electricity usage from December 2006 to November 2010. We chose this domain as a representative time-series forecasting task with connection to everyday decision-making: patterns are intuitive (e.g., higher consumption in winter due to heating) and LLMs can draw on common knowledge to produce meaningful explanations.
@@ 3.5 Prompting Strategies
We select eight prominent strategies from the literature (sahoo2024systematic), spanning simple to complex approaches: zero-shot and few-shot (direct instruction with and without examples), chain-of-thought (zero- and few-shot variants, excluding DeepSeek-R1 which natively uses CoT), meta-prompting (structured 6-step template), reflexion (iterative self-critique), role-based (expert persona), and self-consistency (ensemble of multiple outputs). These range from single-pass methods to multi-iteration approaches. All prompts share a common structure: a system message establishing the task (help a non-technical user understand the forecast), and a human message providing context including model performance metrics, the predicted value, instance features, and XAI attributions when applicable. See Appendix B for complete templates.
We constrain explanations to $\leq$200 words and $\leq$6 bullet points, reflecting realistic user attention spans based on cognitive load research (mark2023attention; brysbaert2019reading) and controlling for verbosity bias (Section 2.4).
@@ 3.6 LLMs for NLE Generation
We use three LLMs spanning different scales and architectures: Llama-3-8B (grattafiori2024llama), a small open-source model run locally; GPT-4o (openai2024gpt4o), a large commercial model; and DeepSeek-R1 (guo2025deepseekr1), a 671B-parameter Mixture-of-Experts model with 37B active parameters, optimized for reasoning via reinforcement learning. This selection examines how model scale and reasoning specialization affect explanation quality. We use temperature $\tau=1.0$ for all strategies to encourage diverse outputs, except self-consistency’s synthesis step ($\tau=0.2$) which requires deterministic aggregation of multiple drafts into a single cohesive explanation (wang2023selfconsistency).
@@ 3.7 Evaluation Framework
We employ G-Eval (liu2023geval) for pointwise scoring with two judge models: GPT-4 and DeepSeek-R1. We exclude Llama as a judge due to its smaller parameter count – research shows that judge capability scales with model size, and weaker models produce less reliable evaluations (gu2024llmasjudge). Each dimension uses chain-of-thought evaluation steps followed by probability-weighted scoring (see Section 2.4). We define four evaluation dimensions grounded in XAI evaluation literature (hoffman2018metrics; miller2019explanation):
@@
  • Lay User Relevancy: Understandability for non-technical users. miller2019explanation notes that “practitioner satisfaction is fundamental to effective explanatory strategies.”
  • Expert Relevancy: Technical soundness for ML experts. ribeiro2016why assert that “good explanations should provide additional insights” enabling model assessment.
  • Explanation Helpfulness: Overall pragmatic usefulness. miller2019explanation notes individuals evaluate explanations “according to pragmatic usefulness.”
Each rubric uses a 1–5 scale with detailed criteria (see Appendix D). Following best practices from gu2024llmasjudge, scores are averaged across both judges to mitigate self-preference bias.
@@ 4 Experiments and Results
For each research question, we test overall effects using one-way ANOVA. Beyond $p$-values, we report omega-squared ($\omega^{2}$) (kroes2023demystifying) as a measure of effect size – the proportion of variance in scores attributable to the factor (e.g., $\omega^{2}=.06$ for XAI means 6% of score variation is due to XAI choice). Following field2013discovering, we interpret: <.01<.01 negligible, .01–.06 small, .06–.14 medium, >.14>.14 large. Where ANOVA is significant, we conduct pairwise $t$-tests reporting Cohen’s $d$ (cohen1988statistical) – the standardized difference between group means. We interpret: <.2<.2 negligible, .2–.5 small, .5–.8 medium, >.8>.8 large.
Because we test five evaluation dimensions simultaneously, we apply Benjamini-Hochberg FDR correction (benjamini1995controlling) to control false discovery rates, reporting both raw $p$ and corrected $p_{\text{FDR}}$ values. We also fit factorial models (maxwell2017designing) controlling for other experimental factors (e.g., testing XAI effect while controlling for LLM and Strategy), and test for interactions to check whether effects vary across conditions.
@@ 5.1 What the Evaluation Dimensions Reveal
Lay Relevancy captures whether non-experts can understand the prediction rationale, assess trustworthiness, and derive actionable insights. It shows no XAI benefit but responds to ML model quality ($\omega^{2}=.12$) and prompting strategy ($\omega^{2}=.12$). The absent XAI effect aligns with research showing that lay users struggle to interpret feature attributions (salih2025perspective; alufaisan2020explainable). Better-performing models provide richer context for explanations; how the LLM frames that context matters equally.
@@
SHAP and LIME show no significant differences across any dimension. While these methods differ algorithmically (SHAP uses Shapley values from game theory, LIME fits local surrogate models), both produce similar feature attribution outputs for tabular data (salih2025perspective). The bottleneck for NLE quality appears to be the LLM’s translation of attributions into narrative, not the XAI method itself.
@@ 6 Conclusion
Evaluation. We rely on LLM-based evaluation (G-Eval) rather than human annotators. Despite our bias mitigation strategies, LLM judges may have systematic biases not captured by our design. More importantly, we measure explanation quality but not utility. Research shows that explanation quality and user performance are often weakly correlated (schemmer2022should): users may over-rely on confident-sounding explanations or ignore technically accurate but complex ones. DeepSeek-R1 produces higher-quality explanations than GPT-4o, but whether users make better decisions with them remains untested. Similarly, the interpretability paradox concerns NLE quality, not whether users actually understand SARIMAX better when explained by humans versus LLMs.
@@ A.1 Data Preprocessing
We use the UCI Individual Household Electric Power Consumption dataset (hebrail2012individual), containing 2,075,259 minute-level measurements from a household in Sceaux, France (near Paris), recorded from December 2006 to November 2010. The preprocessing pipeline: (1) handles missing values ($\approx$1.25%) via linear interpolation, (2) converts kilowatt-minutes to kilowatt-hours, and (3) aggregates to weekly totals anchored on Mondays (ISO week alignment):
@@ Appendix E Statistical Analysis Details
    One-way ANOVA with omega-squared ($\omega^{2}$) effect sizes (kroes2023demystifying). We report $F$-statistics, raw $p$-values, FDR-corrected $p$-values (benjamini1995controlling), and effect size classifications.
@@
    Pairwise comparisons using independent $t$-tests with Cohen’s $d$ effect sizes (cohen1988statistical) and Benjamini-Hochberg FDR correction applied within each dimension.
@@
    Welch ANOVA as robustness check (welch1951comparison), which does not assume equal variances across groups – important given our unbalanced designs (e.g., CoT strategies exclude DeepSeek-R1).
@@
    Games-Howell post-hoc tests (games1976pairwise) for significant dimensions, which are robust to unequal variances and sample sizes.
@@
Following cohen1988statistical and kroes2023demystifying:
@@
We apply Benjamini-Hochberg FDR correction (benjamini1995controlling) to control false discovery rate at $\alpha=.05$. FDR correction is applied: (1) across dimensions within each ANOVA, and (2) within each dimension for pairwise comparisons.

### v2

@@ Abstract
Fabian Lukassen1, Jan Herrmann2, Christoph Weisser4, Benjamin Saefken1,3, Thomas Kneib1 1University of Göttingen  2BASF SE  3TU Clausthal  4 Hochschule Bielefeld fabian.lukassen@stud.uni-goettingen.de, jan.herrmann@basf.com, christoph.weisser@hsbi.de benjamin.saefken@tu-clausthal.de, tkneib@uni-goettingen.de
@@ 1 Introduction
The deployment of ML models in critical domains such as healthcare, finance, and energy management has intensified the need for model interpretability (Guidotti et al., 2018). XAI methods like SHAP (Lundberg and Lee, 2017) and LIME (Ribeiro et al., 2016) have emerged as essential tools, providing quantitative feature attributions that reveal which inputs drive model predictions. However, these technical outputs – typically numerical importance scores or attribution values – remain inaccessible to non-expert users who lack the statistical background to interpret them (Zytek et al., 2024a).
@@
Recent work has demonstrated that LLMs can transform XAI output into NLEs that non-experts can better understand (Figure 1), with user studies validating this approach (Zytek et al., 2024b; Cedro and Martens, 2024).
@@
However, previous work on transforming XAI outputs into NLEs does not systematically study the effects of ML model choice, XAI method, LLM, and prompting strategy on NLE quality. We address this gap with a factorial study grounded in Design of Experiments (DOE) principles (Collins et al., 2014), focusing on time-series forecasting. A domain where temporal dependencies and seasonal patterns make XAI outputs especially challenging to interpret. The factorial combination of these four factors (see Figure 2 for their levels) yields 660 NLEs, which we evaluate using G-Eval (Liu et al., 2023), an LLM-as-judge framework with dual judges and four evaluation criteria. Our main contributions are:
@@ 2.1 Explainable AI Methods
Among post-hoc explanation methods, SHAP and LIME have emerged as the two most widely adopted techniques for interpreting black-box models (Garreau and von Luxburg, 2020).
SHAP (Lundberg and Lee, 2017) provides theoretically grounded attributions based on Shapley values from cooperative game theory. For a prediction $f(x)$ with feature set $\mathcal{F}$, SHAP assigns each feature $i$ an importance value $\phi_{i}$ computed as a weighted average of marginal contributions across all possible feature coalitions $S$:
@@
LIME (Ribeiro et al., 2016) takes a different approach, fitting interpretable local surrogate models. Given instance $x$ and black-box model $f$, LIME learns a locally faithful linear model $g$ by solving:
@@
Both methods have been widely adopted (Guidotti et al., 2018), applied in complex machine learning pipelines (Semnani et al., 2024, 2025), and extended to specialized variants including TreeSHAP (Lundberg et al., 2020). However, despite their mathematical rigor, XAI outputs require statistical expertise to interpret, particularly for time-series data where temporal dependencies add complexity (Theissler et al., 2022). Zytek et al. (2024a) identify this as a fundamental barrier, noting that “feature importance scores are only useful if users can understand what they mean.”
In contrast, classical statistical models like SARIMAX offer inherent interpretability through coefficients and confidence intervals (Box et al., 2015).
@@ 2.2 Natural Language Explanations from XAI
The transformation of XAI outputs to natural language has progressed from template-based systems to LLM-based approaches. Martens et al. (2025) introduced XAIstories, generating narratives from SHAP values, with user studies showing participants found them convincing and useful for communicating with laypeople. Zytek et al. (2024b) developed Explingo, also using SHAP, combining XAI-to-narrative transformation with an LLM-based grader. Cedro and Martens (2024) extended this to graph neural networks using GNNExplainer. For time-series specifically, Shajalal et al. (2024) used DeepLIFT for LSTM-based energy forecasting, while Aksu et al. (2024) proposed XForecast with simulatability metrics, finding that an LLM’s numerical reasoning ability matters more than model size for accurate explanations. Domain applications include chemistry (Wellawatte and Schwaller, 2024).
@@ 2.3 LLM Prompting Strategies
The effectiveness of LLMs depends heavily on prompting. Foundational techniques include zero-shot and few-shot in-context learning (Brown et al., 2020), while more advanced strategies target reasoning and reliability. Chain-of-thought (CoT) prompting elicits step-by-step reasoning (Wei et al., 2022), which self-consistency extends by sampling multiple reasoning paths and selecting the most common answer (Wang et al., 2023). Other approaches include reflexion, where models iteratively critique and refine their outputs (Shinn et al., 2023); meta-prompting, which uses structured templates for multi-step reasoning (Suzgun and Kalai, 2024); and role-based prompting, which assigns personas to guide generation (Kong et al., 2024).
@@ 2.4 LLM-as-Judge Evaluation
Traditional metrics (BLEU, ROUGE) correlate poorly with human judgment for open-ended text generation (Liu et al., 2023). The LLM-as-judge paradigm has emerged as a scalable alternative (Zheng et al., 2023), with three main evaluation paradigms (Gu et al., 2024): pointwise (scoring individual outputs independently), pairwise (comparing two outputs to select the better one), and listwise (ranking multiple outputs simultaneously).
Liu et al. (2023) introduced G-Eval, a pointwise framework combining chain-of-thought prompting with probability-weighted scoring. G-Eval first generates evaluation steps from a criterion (e.g., “coherence”), then uses these steps to assess outputs via a form-filling paradigm. To obtain fine-grained scores, G-Eval computes:
@@
However, LLM judges exhibit systematic biases (Gu et al., 2024): positional bias (preference for first or last positions in ranking tasks), self-preference bias (favoring outputs from the same LLM family), and verbosity bias (favoring longer responses). Mitigation strategies include randomizing presentation order, using judges from different model families than generators, and supplementing absolute scores with ranking-based evaluation.
@@ 3.2 Time-Series Forecasting Task
We use the UCI Individual Household Electric Power Consumption dataset (Hébrail and Bérard, 2012), containing 2,075,259 minute-level measurements of a French household’s electricity usage from December 2006 to November 2010. We chose this domain as a representative time-series forecasting task with connection to everyday decision-making: patterns are intuitive (e.g., higher consumption in winter due to heating) and LLMs can draw on common knowledge to produce meaningful explanations.
@@ 3.5 Prompting Strategies
We select eight prominent strategies from the literature (Sahoo et al., 2024), spanning simple to complex approaches: zero-shot and few-shot (direct instruction with and without examples), chain-of-thought (zero- and few-shot variants, excluding DeepSeek-R1 which natively uses CoT), meta-prompting (structured 6-step template), reflexion (iterative self-critique), role-based (expert persona), and self-consistency (ensemble of multiple outputs). These range from single-pass methods to multi-iteration approaches. All prompts share a common structure: a system message establishing the task (help a non-technical user understand the forecast), and a human message providing context including model performance metrics, the predicted value, instance features, and XAI attributions when applicable. See Appendix B for complete templates.
We constrain explanations to $\leq$200 words and $\leq$6 bullet points, reflecting realistic user attention spans based on cognitive load research (Mark, 2023; Brysbaert, 2019) and controlling for verbosity bias (Section 2.4).
@@ 3.6 LLMs for NLE Generation
We use three LLMs spanning different scales and architectures: Llama-3-8B (Grattafiori and others, 2024), a small open-source model run locally; GPT-4o (OpenAI, 2024), a large commercial model; and DeepSeek-R1 (Guo et al., 2025), a 671B-parameter Mixture-of-Experts model with 37B active parameters, optimized for reasoning via reinforcement learning. This selection examines how model scale and reasoning specialization affect explanation quality. We use temperature $\tau=1.0$ for all strategies to encourage diverse outputs, except self-consistency’s synthesis step ($\tau=0.2$) which requires deterministic aggregation of multiple drafts into a single cohesive explanation (Wang et al., 2023).
@@ 3.7 Evaluation Framework
We employ G-Eval (Liu et al., 2023) for pointwise scoring with two judge models: GPT-4 and DeepSeek-R1. We exclude Llama as a judge due to its smaller parameter count – research shows that judge capability scales with model size, and weaker models produce less reliable evaluations (Gu et al., 2024). Each dimension uses chain-of-thought evaluation steps followed by probability-weighted scoring (see Section 2.4). We define four evaluation dimensions grounded in XAI evaluation literature (Hoffman et al., 2018; Miller, 2019):
@@
  • Lay User Relevancy: Understandability for non-technical users. Miller (2019) notes that “practitioner satisfaction is fundamental to effective explanatory strategies.”
  • Expert Relevancy: Technical soundness for ML experts. Ribeiro et al. (2016) assert that “good explanations should provide additional insights” enabling model assessment.
  • Explanation Helpfulness: Overall pragmatic usefulness. Miller (2019) notes individuals evaluate explanations “according to pragmatic usefulness.”
Each rubric uses a 1–5 scale with detailed criteria (see Appendix D). Following best practices from Gu et al. (2024), scores are averaged across both judges to mitigate self-preference bias.
@@ 4 Experiments and Results
For each research question, we test overall effects using one-way ANOVA. Beyond $p$-values, we report omega-squared ($\omega^{2}$) (Kroes and Finley, 2023) as a measure of effect size – the proportion of variance in scores attributable to the factor (e.g., $\omega^{2}=.06$ for XAI means 6% of score variation is due to XAI choice). Following Field (2013), we interpret: <.01<.01 negligible, .01–.06 small, .06–.14 medium, >.14>.14 large. Where ANOVA is significant, we conduct pairwise $t$-tests reporting Cohen’s $d$ (Cohen, 1988) – the standardized difference between group means. We interpret: <.2<.2 negligible, .2–.5 small, .5–.8 medium, >.8>.8 large.
Because we test five evaluation dimensions simultaneously, we apply Benjamini-Hochberg FDR correction (Benjamini and Hochberg, 1995) to control false discovery rates, reporting both raw $p$ and corrected $p_{\text{FDR}}$ values. We also fit factorial models (Maxwell et al., 2017) controlling for other experimental factors (e.g., testing XAI effect while controlling for LLM and Strategy), and test for interactions to check whether effects vary across conditions.
@@ 5.1 What the Evaluation Dimensions Reveal
Lay Relevancy captures whether non-experts can understand the prediction rationale, assess trustworthiness, and derive actionable insights. It shows no XAI benefit but responds to ML model quality ($\omega^{2}=.12$) and prompting strategy ($\omega^{2}=.12$). The absent XAI effect aligns with research showing that lay users struggle to interpret feature attributions (Salih et al., 2025; Alufaisan et al., 2021). Better-performing models provide richer context for explanations; how the LLM frames that context matters equally.
@@
SHAP and LIME show no significant differences across any dimension. While these methods differ algorithmically (SHAP uses Shapley values from game theory, LIME fits local surrogate models), both produce similar feature attribution outputs for tabular data (Salih et al., 2025). The bottleneck for NLE quality appears to be the LLM’s translation of attributions into narrative, not the XAI method itself.
@@ 6 Conclusion
Evaluation. We rely on LLM-based evaluation (G-Eval) rather than human annotators. Despite our bias mitigation strategies, LLM judges may have systematic biases not captured by our design. More importantly, we measure explanation quality but not utility. Research shows that explanation quality and user performance are often weakly correlated (Schemmer et al., 2022): users may over-rely on confident-sounding explanations or ignore technically accurate but complex ones. DeepSeek-R1 produces higher-quality explanations than GPT-4o, but whether users make better decisions with them remains untested. Similarly, the interpretability paradox concerns NLE quality, not whether users actually understand SARIMAX better when explained by humans versus LLMs.
@@ A.1 Data Preprocessing
We use the UCI Individual Household Electric Power Consumption dataset (Hébrail and Bérard, 2012), containing 2,075,259 minute-level measurements from a household in Sceaux, France (near Paris), recorded from December 2006 to November 2010. The preprocessing pipeline: (1) handles missing values ($\approx$1.25%) via linear interpolation, (2) converts kilowatt-minutes to kilowatt-hours, and (3) aggregates to weekly totals anchored on Mondays (ISO week alignment):
@@ Appendix E Statistical Analysis Details
    One-way ANOVA with omega-squared ($\omega^{2}$) effect sizes (Kroes and Finley, 2023). We report $F$-statistics, raw $p$-values, FDR-corrected $p$-values (Benjamini and Hochberg, 1995), and effect size classifications.
@@
    Pairwise comparisons using independent $t$-tests with Cohen’s $d$ effect sizes (Cohen, 1988) and Benjamini-Hochberg FDR correction applied within each dimension.
@@
    Welch ANOVA as robustness check (Welch, 1951), which does not assume equal variances across groups – important given our unbalanced designs (e.g., CoT strategies exclude DeepSeek-R1).
@@
    Games-Howell post-hoc tests (Games and Howell, 1976) for significant dimensions, which are robust to unequal variances and sample sizes.
@@
Following Cohen (1988) and Kroes and Finley (2023):
@@
We apply Benjamini-Hochberg FDR correction (Benjamini and Hochberg, 1995) to control false discovery rate at $\alpha=.05$. FDR correction is applied: (1) across dimensions within each ANOVA, and (2) within each dimension for pairwise comparisons.